# Curating Demonstrations using Online Experience

### v1

@@ IV-B Demonstration Filtering
We can then use the filtered dataset $\mathcal{D}_{\text{demo, filt}}$ to train a new policy, $\pi_{\text{final}}$, either from scratch or by fine-tuning the initial policy which was trained on the full demonstration dataset. In a variant of Demo-SCORE requiring no additional supervision, we can apply our same trained classifier on the successful rollouts themselves, obtaining $\mathcal{D}_{\text{rollouts, filt}}$, and we can then train on the union of $\mathcal{D}_{\text{demo, filt}}$ and $\mathcal{D}_{\text{rollouts, filt}}$ in this last stage. By filtering out unreliable demonstrations, Demo-SCORE helps ensure that the policy learns the most effective strategies, ultimately improving the performance and robustness of the learned robot behavior. Our full method is summarized in Algorithm 1.
@@
6:  $\mathcal{D}_{\text{demo, filt}}\leftarrow\{\tau_{j}\mid\bar{q}_{{\phi^{*}}}(\tau_{j})>\gamma,\tau_{j}\in\mathcal{D}_{\text{demo}}\}$
7:  $\pi_{\text{final}}\leftarrow$ Use $\mathcal{D}_{\text{demo, filt}}$ to re-train or fine-tune $\pi$
@@ Results
TABLE I: Subtask Percent Success in Jellybean Task. We report the percent success rate for each subtask in the Jellybean task, the aggregated score (the percentage of all subtasks completed), and the percent success rate on the full Jellybean task. On this challenging task, Demo-SCORE is able to curate the data so that the robot is able to even achieve success on the full task.
@@ V-D Variations of Demo-SCORE
TABLE III: Ablation on number of rollouts needed. Demo-SCORE achieves strong performance using as few as 10 rollouts per checkpoint for classifier training.
In this ablation, we test whether using fewer rollouts, such as 50 or even 10, from the first $C-1$ checkpoints to train the classifier, and only 50 rollouts from the last checkpoint for the validation set, leads to the similar performance for Demo-SCORE. This would be especially appealing for real-world settings where collecting rollouts may be expensive or time-consuming. In Table III, we find that using fewer rollouts leads to only a small decrease in performance, suggesting that only a small number of rollouts are needed to train an effective classifier. We note that other methods that leverage online rollouts have a much larger decrease in performance with fewer rollouts. Notably, with 10 rollouts per checkpoint and fewer than 100 rollouts total, Demo-SCORE achieves 92% average success rate across data mixtures, over 10% higher than the next best method, Auto-IL.
@@ A-D Additional Simulated Result Tables
In Table IV, we report the aggregated results from Figure 4. Demo-SCORE demonstrates 15-35% improvement in average success rate across data mixtures over the base policy and roughly 2x lower failure rate compared to the next best method using online rollouts in both evaluation suites. We provide the number of demonstrations filtered out by Demo-SCORE in Table V for the square peg task, showing that Demo-SCORE adjusts the amount of demos filtered depending on the quality mixture of the original demonstrations. In Table, VI, we show that the performance of DemoSCORE on the square task is consistent as we vary the number of hidden units and layers in the step classifier MLP.
@@ A-E Additional Details on Jellybean Task
Base Policy Training Loss RCP Auto-IL Demo-SCORE
Peg insertion (100 rollouts) 58.3 71.3 61.3 79.9 90.8
Square peg (100 rollouts) 79.9 79.2 87.1 88.2 94.4
TABLE IV: Average success rate across data mixtures. In this table, we aggregate results from Figure 4.
Num Demos Filtered Out 104 93 35 142 200 400
TABLE V: Number of demonstrations filtered by Demo-SCORE. Demo-SCORE is effective even with strongly lopsided heterogeneous mixtures and adjusts accordingly depending on how much should be filtered from a given dataset.
Success Rate 79.9 94.4 95.8 95.3 94.4 94.0 95.1
TABLE VI: Ablation on Classifier Size. For the square task, we report the average percent success rate for the across data splits when using Demo-SCORE with step MLP classifiers of different sizes. For each MLP architecture, we report the number of hidden units in each hidden MLP layer and the average success rate of the final Demo-SCORE policies. The main experiments used an “[8, 8]" architecture with two hidden layers with 8 units each.

### v2

@@ IV-B Demonstration Filtering
We can then use the filtered dataset $\mathcal{D}_{\text{demo, filt}}$ to train a new policy, $\pi_{\text{final}}$, either from scratch or by fine-tuning the initial policy which was trained on the full demonstration dataset. We then apply our same trained classifier on the successful rollouts themselves, obtaining $\mathcal{D}_{\text{rollouts, filt}}$, and we can then train on the union of $\mathcal{D}_{\text{demo, filt}}$ and $\mathcal{D}_{\text{rollouts, filt}}$ in this last stage. By filtering out unreliable demonstrations, Demo-SCORE helps ensure that the policy learns the most effective strategies, ultimately improving the performance and robustness of the learned robot behavior. Our full method is summarized in Algorithm 1.
@@
6:  $\mathcal{D}_{\text{demo, filt}},\mathcal{D}_{\text{rollouts, filt}}\leftarrow\{\tau_{j}\mid\bar{q}_{{\phi^{*}}}(\tau_{j})>\gamma,\tau_{j}\in\mathcal{D}_{\text{demo}}\cup\mathcal{D}_{\text{rollouts}}\}$
7:  $\pi_{\text{final}}\leftarrow$ Use $\mathcal{D}_{\text{demo, filt}}\cup\mathcal{D}_{\text{rollouts, filt}}$ to re-train & fine-tune $\pi$
@@ Results
TABLE I: We report the percent success rate for each subtask in the Jellybean task, the aggregated score (the percentage of all subtasks completed), and the percent success rate on the full Jellybean task. On this challenging task, Demo-SCORE is able to curate the data so that the robot is able to even achieve success on the full task.
@@ V-D Variations of Demo-SCORE
Peg insertion (100 rollouts) 58.9 71.3 61.3 79.9 90.8
Peg insertion (50 rollouts) 58.9 71.3 59.6 74.4 88.0
Peg insertion (25 rollouts) 58.9 71.3 56.3 70.2 89.1
Peg insertion (10 rollouts) 58.9 71.3 58.9 67.8 87.7
TABLE III: Ablation on number of rollouts needed. On two simulated tasks, Demo-SCORE achieves strong performance using as few as 10 rollouts per checkpoint for classifier training.
In this ablation, we test whether using fewer rollouts, such as 50 or even 10, from the first $C-1$ checkpoints to train the classifier, and only 50 rollouts from the last checkpoint for the validation set, leads to the similar performance for Demo-SCORE. This would be especially appealing for real-world settings where collecting rollouts may be expensive or time-consuming. In Table III, we find that using fewer rollouts leads to only a small decrease in performance, suggesting that only a small number of rollouts are needed to train an effective classifier. We note that other methods that leverage online rollouts have a much larger decrease in performance with fewer rollouts. Notably, with 10 rollouts per checkpoint and fewer than 100 rollouts total, Demo-SCORE achieves 92% average success rate across data mixtures, over 10% higher than the next best method, Auto-IL, on the square peg task. Similarly, with only 10 rollouts per training checkpoint and 80 total rollouts, Demo-SCORE achieves an 87.7 % success rate, compared to a 58.9% success rate for the base policy.
@@ Appendix A Appendix
Num Demos Filtered Out 104 93 35 142 200 400
TABLE IV: Number of demonstrations filtered by Demo-SCORE. Demo-SCORE is effective even with strongly lopsided heterogeneous mixtures and adjusts accordingly depending on how much should be filtered from a given dataset.
Success Rate 79.9 94.4 95.8 95.3 94.4 94.0 95.1
TABLE V: Ablation on Classifier Size. For the square task, we report the average percent success rate for the across data splits when using Demo-SCORE with step MLP classifiers of different sizes. For each MLP architecture, we report the number of hidden units in each hidden MLP layer and the average success rate of the final Demo-SCORE policies. The main experiments used an “[8, 8]" architecture with two hidden layers with 8 units each.
@@ A-D Additional Simulated Result Tables
We provide the number of demonstrations filtered out by Demo-SCORE in Table IV for the square peg task, showing that Demo-SCORE adjusts the amount of demos filtered depending on the quality mixture of the original demonstrations. In Table, V, we show that the performance of DemoSCORE on the square task is consistent as we vary the number of hidden units and layers in the step classifier MLP.
@@ A-E Additional Details on Jellybean Task
A-F Impact of Demo-SCORE Filtering on OOD Performance
One potential concern is that Demo-SCORE narrows the state distribution in the original dataset, which may affect generalization capabilities. To understand this, we evaluate two generalization settings in the square peg domain: (1) OOD initial conditions, and (2) different train-test splits across in-task variations of different initial conditions.
For (1), in Table VI, we show the success rate of the square task policies trained on each data split within a modified environment where the xy dimensions of the region where the square nut is initialized are expanded by 50% (easy) and 100% (hard) beyond the standard environment. As the demos used to train these policies were all from the original environment, the nut’s initial position is OOD in these expanded regions relative to the datasets used to train these policies. We find that Demo-SCORE (with either full episode filtering or chunk-level filtering) does not harm OOD generalization in this setting compared to the base policy or Auto-IL and even improves performance whe the environment is initialized in these OOD states.
For (2), our goal is to test the test the policy’s generalization ability to initial conditions that are both in-distribution and OOD but have mostly or only suboptimal data in some regions. Specifically, we train new policies on new datasets in which there is a correlation between the initial state of the nut and the strategy used in each demo. We divide the initial state space of the nut by the orientation of the nut, defining "even" and "odd" regions corresponding to the quadrant in which the angle of rotation around the z-axis of the nut falls. (In other words, if the angle by which the nut is rotated around the z-axis relative to its reference position falls between 0 and 90 degrees or 180 and 270 degrees, it is considered within the "odd" quadrant region).
In these new datasets, we control how many Human and Square A demos there are in which the nut’s initial orientation is Odd or Even. Each dataset has 200 total demonstrations, 100 for the odd and even region. In the Even region, we vary the percentage of demos that are on the Square A mode from 100% to 50%, in increments of 10%. We either chose the opposite ratio of modes in the Odd region or keep 50 demos from each modes in the odd region. We present the success rates of the policies trained on these datasets in Table VII.
Demo-SCORE improves performance when different ratios of high-quality and suboptimal demonstrations exist for different initial conditions and does not degrade performance for states dominated by suboptimal demonstrations, suggesting that Demo-SCORE does not filter out suboptimal data that contributes to generalization success. However, in some of the most extreme cases (such as one data split where 100% of demos in the even region of initial state space have the Square A strategy), Demo-SCORE can slightly under-perform other methods. Nonetheless, these results suggest that Demo-SCORE generally does not excessively narrow the state distribution in a way that harms policy generalization. Intuitively, Demo-SCORE aims to keep any data that contributes to policy success and therefore may retain suboptimal strategies if they appear in successful rollouts. Additionally, if there are settings where filtering may be overly restrictive, this can be mitigated by common alternatives, e.g. reweighting. Our key contribution is an approach for estimating the quality of a demo or chunk, and this score can be applied in various ways, including but not limited to filtering.
50% Expanded XY 100% Expanded XY
@@ A-F Impact of Demo-SCORE Filtering on OOD Performance
Data Split Base Policy Auto-IL Demo-SCORE Demo-SCORE (chunk) Base Policy Auto-IL Demo-SCORE Demo-SCORE (chunk)
100 Human 100 SquareA 65.2 71.9 85.9 85.9 51.6 63.3 73 72.3
100 Human, 50/50 Square A/B 66.4 77.7 88.7 85.2 54.3 64.8 74.2 72.3
200 Human, 20 Square A 89.8 90.6 92.6 92.2 80.1 80.9 83.2 82
200 Human, 100 Square A 73.4 84.8 89.1 90.6 64.5 75.4 81.3 82
200 Human, 200 Square A 75.4 87.1 93 94.5 61.7 82.4 83.4 80.5
200 Human, 400 Square A 68 76.2 94.1 94.9 54.7 65.6 83.6 84.4
Average 73.0 81.4 90.6 90.6 61.2 72.2 79.8 78.9
TABLE VI: OOD Evaluation of Square Task Policies. The success rates of the base policies, policies trained with Auto-IL, and two variants of Demo-SCORE, on OOD environments. The xy dimensions of the region in which the square nut’s initial position is sampled (uniformly) are expanded by 50% and 100% relative to the original environment. The demonstrations used to train the evaluated policies come from the original environment, and these policies are thus OOD on the modified environments. Demo-SCORE outperforms the base policy and the Auto-IL policy when the environment is initialized to OOD states.
Standard Environment 50% Expanded XY 100% Expanded XY
Data Mixture Base Policy Auto-IL Demo-SCORE Base Policy Auto-IL Demo-SCORE Base Policy Auto-IL Demo-SCORE
100% SqA Even, 100% Human Odd 54.3 62.9 69.9 57.0 56.2 63.7 46.9 49.6 51.2
90% SqA Even, 90% Human Odd 59.8 63.3 78.1 57.4 60.5 71.5 50.4 52.3 62.9
80% SqA Even, 80% Human Odd 66.4 73.4 86.7 62.9 68.4 80.9 53.9 56.6 68.4
70% SqA Even, 70% Human Odd 69.5 83.6 91.4 68.8 75.4 89.8 62.5 60.5 78.1
60% SqA Even, 60% Human Odd 67.2 78.1 88.7 64.8 77.0 86.3 57.0 67.6 70.7
50% SqA Even, 50% Human Odd 67.6 76.2 91.4 65.2 71.9 85.5 57.0 62.1 75.4
100% SqA Even, 50% Human Odd 45.7 56.2 52.3 43.8 49.6 51.2 43.8 40.6 39.8
90% SqA Even, 50% Human Odd 48.8 60.5 69.9 44.5 61.3 65.6 41.8 48.4 52.0
80% SqA Even, 50% Human Odd 61.7 68.0 85.2 60.9 64.8 79.7 53.1 53.9 66.4
70% SqA Even, 50% Human Odd 63.3 69.5 78.1 59.8 64.5 75.8 53.1 58.6 64.1
60% SqA Even, 50% Human Odd 67.6 76.2 89.8 65.2 77.0 86.3 59.8 69.1 71.1
TABLE VII: Success Rates on Correlated Datasets. The success rates on the square task of base policies, Auto-IL, and Demo-SCORE policies when the original data mixture has a correlation between the initial orientation of the square nut and the strategy exhibited in each demo as described in section A-F. Each row corresponds to an initial data mixture and is denoted by the percentage of demos within the even region that are employ the Square A strategy and percentage of demos within the odd region that are Human demos (which employ the same strategy). We present evaluations for the standard environment, as well as OOD evaluations on a modified environments where the square nut’s initial xy position is initialized within a region expanded 50 or 100% relative the standard environment. Demo-SCORE generally outperforms the base policy and Auto-IL even when there is a correlation between initial state and demonstrated strategy and the evaluation environment is OOD.